# Practical Use Cases of Neutral Atoms Quantum Computers

### v1

@@ A.2 Digital computing mode
Controlled-NOT (or CX) can be implemented by applying two Hadamard gates interleaved by a CZ gate (see box 1 in [9]). The CZ is implemented with three consecutive laser pulses, two of them acting on the control (1 and 3, box 1 in [9]) and one on the target (number 2, box 1 in [9]). The idea is that when the starting state is either $|00\rangle,|01\rangle$ or $|10\rangle$ the whole two atoms system gets a phase equal to $e^{i\pi}=-1$ due to either the $2\pi$ pulse $2$ or the combination of $\pi$ pulses 1 and 3. Notice that if the initial state is $|00\rangle$, pulse 2 becomes off-resonant due to the shift of the Rydberg level on the target qubit. In the $|11\rangle$ case, all the pulses are off-resonant and no phase is acquired. Therefore, this protocol allows to apply the unitary transformation $diag(-1,-1,-1,1)=e^{i\pi}CZ$ where the global phase $e^{i\pi}$ can be neglected.
@@ A.5 Noise sources in neutral atoms devices
The main noise sources are shown in fig. 1c of [71].
The primary source of errors is due to Black-Body Radiation (BBR). When this occurs, an excited atom "jumps" to a nearby Rydberg level compatibly with selection rules. Since this level cannot be de-excited to $|0\rangle$ to which it is not coupled and cannot decay to the ground state within the time of applications of subsequent gates, this entails a continuous blockade effect on nearby atoms, potentially resulting in correlated $Z$-type errors. The overall effect of BBR-induced errors can be suppressed by increasing the principal quantum number $n$ of Rydberg levels or decreasing the temperature since the rate of BBR transition is proportional to $T$ (see eq. (3) in [71]).
When a spontaneous emission event occurs in an excited atoms, we talk about Radiative Decay (RD). The atom in the Rydberg state decays into one of the levels in the manifold containing the $|1\rangle$ state (aqua green fig. 1c of [71]). This entails $Z$-type correlated errors if the RD event occur during entangling gates. Furthermore, RD rate is temperature-independent and decreases with $n$. Hence, RD dominates over BBR in the low $T$ and small $n$ limit [71]. The RD error class includes a peculiar subclass containing the so called Intermediate state scattering, consisting in decays from intermediate state, e.g. $5P_{3/2}$ (gray, fig. 1c of [71]) to three possible ground states, including $|1\rangle$. This error can be suppressed by increasing laser detuning and laser power [71].
@@ B.1 The Quantum Approximate Optimization Algorithm
The Quantum Approximate Optimization Algorithm, a.k.a. QAOA, has been introduced by Farhi et al. in 2014 [36]. The core idea is to drive the system evolution with two non-commuting Hamiltonians, called the cost Hamiltonian $H_{C}$, i.e. the energy operator whose ground state encodes the solution of the problem, and the mixer Hamiltonian $H_{M}$. Since an overwhelming majority of the problem Hamiltonians include terms which consist of Pauli $\sigma^{z}$ operators, the typical choice for the mixer is $H_{M}=\sum_{i}\sigma^{x}_{i}$, because it does not commute with $\sigma^{z}$, i.e. $\sigma^{x}_{i},\sigma^{z}_{j}]\neq 0$. If the cost Hamiltonian itself consists of two non-commuting terms, the evolution operators to be applied will be driven by these two terms. This is the case when one tries to find the ground state of physical models like the Ising model in transverse field. The QAOA imposes to set the initial state of the system to the balanced superposition of all basis states $|+\rangle^{\otimes N}$ where $|+\rangle=\frac{|0\rangle+|1\rangle}{\sqrt{2}}$ is the eigenstate of $\sigma^{x}$ with eigenvalue 1. Parameterized evolution operators driven by the cost and the mixer Hamiltonians are then applied in alternated fashion for $p$ times, where $p$ is called the "depth" of the circuit, a.k.a. its number of layers. At the end of the circuit the energy, i.e. the cost value, is estimated from repeated measurements as $\frac{1}{N}\sum_{i}\langle\psi_{i}|H_{C}|\psi_{i}\rangle$, where $|\psi_{i}\rangle$ is the outcome bitstring of the $i$-th measure and $N$ is the number of measurements. The parameters are then updated by a classical optimizer and the procedure is repeated until the optimizer converges to the best circuit parameters.

### v2

@@ 9 Conclusions
These systems can already provide concrete value for a class of near-term problems despite current hardware limitations, such as restricted coherence times and lack of local addressing. Medium-scale instances of Maximum Independent Set and related graph problems are native to analog encoding in the Rydberg Hamiltonian, as are QUBO models with moderate connectivity requirements. Other near-term applications include quantum-enhanced graph machine learning tasks, where geometric flexibility of atomic arrays enables expressive quantum feature maps.
This study was funded by the European Union - NextGenerationEU, Mission 4, Component 2, Investment 1.4 in the framework of the CN HPC - “National Centre for HPC, Big Data and Quantum Computing”, (CN HPC - CN00000013 – CUP D56G22000380006). The views and opinions expressed are solely those of the authors and do not necessarily reflect those of the European Union, nor can the European Union be held responsible for them.
E.E. acknowledges financial support from the MUR 2022-PRIN Project “Hybrid algorithms for quantum simulators”. E.E. and M.G. were partially funded by INFN IS-Quantum.
@@ A.2 Digital computing mode
Figure 6: Pulse sequence required to implement CZ gate on neutral atoms in digital mode: a $\pi$ pulse is applied to the control qubit $q_{c}$, followed by a $2\pi$ pulse on the target qubit $q_{t}$ and another $\pi$ pulse to de-excite $q_{c}$. If the initial state is either $|00\rangle$, $|01\rangle$ or $|10\rangle$, the system acquires a phase 1-1 due to the successful application of the $2\pi$ pulse of step 2 or the combination of pulse 1 and 3. The system state is left invariant only when the initial state is $|11\rangle$ since all the pulses are off-resonant
@@
Controlled-NOT (or CX) can be implemented by applying two Hadamard gates interleaved by a CZ gate. The CZ is implemented with three consecutive laser pulses, two of them acting on the control and one on the target (see Fig. 6). The idea is that when the starting state is either $|00\rangle,|01\rangle$ or $|10\rangle$ the whole two atoms system gets a phase equal to $e^{i\pi}=-1$ due to either the $2\pi$ pulse $2$ or the combination of $\pi$ pulses 1 and 3. Notice that if the initial state is $|00\rangle$, pulse 2 becomes off-resonant due to the shift of the Rydberg level on the target qubit. In the $|11\rangle$ case, all the pulses are off-resonant and no phase is acquired. Therefore, this protocol allows to apply the unitary transformation $diag(-1,-1,-1,1)=e^{i\pi}CZ$ where the global phase $e^{i\pi}$ can be neglected.
@@ A.5 Noise sources in neutral atoms devices
Figure 7: Main noise sources in neutral atom devices: Black-Body Radiation (BBR, red) consists in the excitation of the system from the Rydberg state $|nS\rangle$ used for computation to a Rydberg $|nP\rangle$ state; Radiative Decay (RD, light blue) consists in the de-excitation of the atom from the Rydberg state to the ground state manifold; Intermediate State Scattering (ISS, purple) consists in the de-excitation of the atom from the intermediate state to the ground state manifold
The main noise sources are shown in Fig. 7.
The primary source of errors is due to Black-Body Radiation (BBR) (red Fig. 7). When this occurs, an excited atom "jumps" to a nearby Rydberg level compatibly with selection rules. Since this level cannot be de-excited to $|0\rangle$ to which it is not coupled and cannot decay to the ground state within the time of applications of subsequent gates, this entails a continuous blockade effect on nearby atoms, potentially resulting in correlated $Z$-type errors. The overall effect of BBR-induced errors can be suppressed by increasing the principal quantum number $n$ of Rydberg levels or decreasing the temperature since the rate of BBR transition is proportional to $T$ and inversely proportional to $n^{2}$ [46].
When a spontaneous emission event occurs in an excited atoms, we talk about Radiative Decay (RD). The atom in the Rydberg state decays into one of the levels in the ground state manifold (light blue Fig. 7). This entails $Z$-type correlated errors if the RD event occur during entangling gates. Furthermore, RD rate is temperature-independent and decreases with $n$. Hence, RD dominates over BBR in the low $T$ and small $n$ limit [46]. The RD error class includes a peculiar subclass containing the so called Intermediate state scattering, consisting in decays from intermediate state, e.g. $6P_{3/2}$ (purple, Fig. 7) to the ground state manifold. This error can be suppressed by increasing laser detuning and laser power [46].
@@ B.1 The Quantum Approximate Optimization Algorithm
Figure 8: Scheme of the Quantum Approximate Optimization Algorithm. The system is prepared in a superposition state via Hadamard gates and alternately evolved with the cost Hamiltonian of the problem $H_{C}$ and a mixing Hamiltonian $H_{M}$ which is chosen such that $H_{C},H_{M}]\neq 0$. The evolution steps are parameterized by angles $\gamma_{1},...,\gamma_{p}$ (for cost layers) and $\beta_{1},...,\beta_{p}$ (for mixing layers). At each iteration the energy of the system is computed and a classical optimizer minimizes the energy by updating the parameter set up to convergence
The Quantum Approximate Optimization Algorithm, a.k.a. QAOA, has been introduced by Farhi et al. in 2014 [67]. Fig. 8 provides a scheme of its workflow. The core idea is to drive the system evolution with two non-commuting Hamiltonians, called the cost Hamiltonian $H_{C}$, i.e. the energy operator whose ground state encodes the solution of the problem, and the mixer Hamiltonian $H_{M}$. Since an overwhelming majority of the problem Hamiltonians include terms which consist of Pauli $\sigma^{z}$ operators, the typical choice for the mixer is $H_{M}=\sum_{i}\sigma^{x}_{i}$, because it does not commute with $\sigma^{z}$, i.e. $\sigma^{x}_{i},\sigma^{z}_{j}]\neq 0$. If the cost Hamiltonian itself consists of two non-commuting terms, the evolution operators to be applied will be driven by these two terms. This is the case when one tries to find the ground state of physical models like the Ising model in transverse field. The QAOA imposes to set the initial state of the system to the balanced superposition of all basis states $|+\rangle^{\otimes N}$ where $|+\rangle=\frac{|0\rangle+|1\rangle}{\sqrt{2}}$ is the eigenstate of $\sigma^{x}$ with eigenvalue 1. Parameterized evolution operators driven by the cost and the mixer Hamiltonians are then applied in alternated fashion for $p$ times, where $p$ is called the "depth" of the circuit, a.k.a. its number of layers. At the end of the circuit the energy, i.e. the cost value, is estimated from repeated measurements as $\frac{1}{N}\sum_{i}\langle\psi_{i}|H_{C}|\psi_{i}\rangle$, where $|\psi_{i}\rangle$ is the outcome bitstring of the $i$-th measure and $N$ is the number of measurements. The parameters are then updated by a classical optimizer and the procedure is repeated until the optimizer converges to the best circuit parameters.